OpenStudy (_excusemyfrench._):

A research study in 2000 found that there were 440 coyotes in a given region. The coyote population is expected to grow at a rate of 17% each year. In what year will the population first exceed 10,000?

OpenStudy (anonymous):

Let t = # of years since 2000 Then we have: y=440(1+0.17)^t

OpenStudy (_excusemyfrench._):

so y=514^2 ??

OpenStudy (anonymous):

do u know how to take ln of both sides?

OpenStudy (anonymous):

10000=440(1+0.17)^t

OpenStudy (anonymous):

\[\frac{ 10000 }{ 440 }=(1+0.17)^t\]

OpenStudy (anonymous):

then we take ln of both sides\[\ln(22.73)=t \ln(1+0.17)\]

OpenStudy (anonymous):

\[\frac{ \ln(22.73) }{ \ln(1+0.17)}=t\]

OpenStudy (anonymous):

what do u get?

OpenStudy (_excusemyfrench._):

22.9?

OpenStudy (anonymous):

r u sure i am getting 19.89

OpenStudy (anonymous):

dont use log use ln

OpenStudy (anonymous):

on ur calculator

OpenStudy (anonymous):

so basically u roud this up and u will get 20 years so 2000+20 which will equal 2020

OpenStudy (anonymous):

so in 2020 the population first exceed 10,000?

OpenStudy (_excusemyfrench._):

um..im confused

OpenStudy (_excusemyfrench._):

the answer is 20?

OpenStudy (anonymous):

no it asks in what year the population will exceed 10000

OpenStudy (_excusemyfrench._):

okay sorry i meant 2020

OpenStudy (anonymous):

yes thats correct

OpenStudy (_excusemyfrench._):

yay omg thank you!!

OpenStudy (_excusemyfrench._):

it was 2019 but thanks for helping me understand it lol

OpenStudy (anonymous):

opps srry i thought cause its coming 19.89 it should be rounded up